4. Show how Charless Law can be derived from the combined gas law.
Answer #1

Combined gas law :- PV/T = k ( constant )

Therefore , P1V1/T1 = P2V2/T2

P1 = Pressure of first gas P2= Pressure of second gas

V1 = Volume of first gas V2=Volume of second gas

T1= Temperature of first gas T2 = Temperature of second gas

In case , the pressure remains constant in P1V1/T1 = P2V2/T2, we can write it as PV1/T1=PV2/T2

Now , cancelling P from both sides ( as it is same ) , We get V1/T1=V2/T2 that is CHARLES' LAW